OpenStudy (anonymous):

Solve the polynomial inequality. Express the solution set in interval notation. x^3-8x^2+6x-48

jhonyy9 (jhonyy9):

x2(x-8)+6(x-8)=(x2+6)(x-8) x2+6=0 x2=-6 x_1,_2=+/- sqrt(-6) x-8=0 x=8

OpenStudy (anonymous):

yes but how do u put it in interval notation

jhonyy9 (jhonyy9):

inequality where is ?

jhonyy9 (jhonyy9):

how is the inequality sign ???

OpenStudy (anonymous):

>

jhonyy9 (jhonyy9):

the polinom need to be geater than 0

OpenStudy (anonymous):

>0

jhonyy9 (jhonyy9):

ok

OpenStudy (anonymous):

i tried to put it in interval notation and ic ant get it

jhonyy9 (jhonyy9):

x=(8,+ infinity) but this is just one part

OpenStudy (anonymous):

whats the other part

jhonyy9 (jhonyy9):

this is all

OpenStudy (anonymous):

okay thanks

jhonyy9 (jhonyy9):

because x2+6 will be always positiv

jhonyy9 (jhonyy9):

my pleasure

OpenStudy (anonymous):

:)